OpenStudy (anonymous):

Write an equation in slope intercept form with the given slope through the given point: slope=8/3, (-2,-7)

OpenStudy (anonymous):

y = mx + c ( equation of a line) m = slope = 8/3 points x = -2 , y = -7 substitute the values -7 = 8/3 * -2 + c -7 = -16/3 + c -7 + 16/3 = c -21 + 16 --------- = c 3 -5/3 = c therefore equation of line y = 8/3x -5/3